# Emilia: A Large-Scale, Extensive, Multilingual, and Diverse Dataset for Speech Generation

### v1

@@ V-C Multilingual and Crosslingual Speech Generation (EQ3)
In this subsection, we conduct experiments to evaluate the effectiveness of the Emilia dataset for training multilingual and crosslingual speech generation models.
@@
For this experiment, we utilize the complete Emilia-Large dataset, which encompasses six languages: English (En), Chinese (Zh), German (De), French (Fr), Japanese (Ja), and Korean (Ko), to train speech generation models. For evaluation, the English (En) evaluation set is derived from Emilia-Test. The Chinese (Zh) evaluation set is randomly sampled from the AISHELL-3 dataset [19], while the evaluation sets for German (De), French (Fr), Japanese (Ja), and Korean (Ko) are sourced from Common Voice [43]. Each evaluation set contains a minimum of 500 reference speech samples. In multilingual experiments, these reference speech samples are directly utilized to synthesize target texts in the same language. For crosslingual experiments, the reference speech samples and the target texts are sampled from evaluation sets of different languages.
@@
The experimental results in Table V demonstrate the effectiveness of the Emilia-Large dataset for multilingual and crosslingual speech generation.
In multilingual generation, where the reference and target languages are identical, both AR+SoundStorm and VoiceBox achieve strong performance across all six languages. Spcifically, AR+SoundStorm exhibits WER ranging from 3.6% (Zh-Zh) to 6.3% (Ko-Ko), with S-SIM between 0.511 (Zh-Zh) and 0.673 (Ko-Ko), and FSD as low as 24.99 (En-En). VoiceBox also shows comparable performance, with WER ranging from 5.2% (De-De) to 6.5% (En-En), S-SIM from 0.557 (Zh-Zh) to 0.683 (De-De), and FSD from 24.34 to 44.71 (Ja-Ja).
In crosslingual generation, where the reference and target languages differ, performance remains robust but experiences moderate degradation. For instance, the WER increases, such as from 6.3% (Ko-Ko) to 8.3% (En-Ko) for AR+SoundStorm. Similarly, the S-SIM decreases, as seen in the comparison between En-En (0.588) and (En-Zh) 0.386 for VoiceBox. The FSD also rises, for example, the FSD for En-Zh in AR+SoundStorm is 99.40 compared to 24.99 for Zh-Zh.
These results highlight two key findings. First, the Emilia-Large dataset facilitates strong multilingual speech generation for both models. Second, crosslingual generation introduces notable challenges: while models retain functionality, performance gaps emerge, particularly for VoiceBox, indicating that the NAR-based architectures may struggle more with cross-language acoustic transfer. Furthermore, when comparing multilingual training with English-only training (Fig. 57), multilingual models exhibit a slight trade-off in English performance. For instance, AR+SoundStorm trained on the multilingual 256k-hour Emilia-Large achieves a WER of 4.9%, an S-SIM of 0.636, and an FSD of 15.31 on the English samples in the evaluation sets. These metrics are slightly worse than those of its monolingual counterpart trained on a 134k-hour English-only dataset (WER=4.5%, S-SIM=0.65, FSD=14.8, as reported in prior results). This suggests that while multilingual data enables crosslingual capabilities, it minimally compromises language-specific performance. This trade-off is an important consideration for practical applications. Future work could explore strategies to narrow these performance gaps.
@@
The Emilia-Large dataset proves highly effective for training multilingual and crosslingual speech generation models. However, we find that crosslingual generation introduces moderate performance degradation, highlighting challenges in cross-language acoustic transfer. These findings underscore the value of the Emilia-Large dataset as a critical resource for advancing multilingual and crosslingual speech generation. Future work could focus on enhancing model adaptability to address crosslingual challenges.
VI Conclusion and Limitation
In conclusion, this work introduces Emilia-Pipe, an effective preprocessing pipeline designed to transform raw in-the-wild speech data into high-quality training datasets for spontaneous and human-like speech generation. Leveraging Emilia-Pipe, we construct Emilia, the largest open-source multilingual speech generation dataset, spanning over 101k hours across six languages, and its extended version, Emilia-Large, encompassing 216k hours. Comparative analyses demonstrate that Emilia significantly outperforms traditional audiobook datasets in generating spontaneous and human-like speech. Our experiments also investigate the relationship between dataset size and speech generation performance, revealing consistent improvements with data scaling, though the trend becomes less pronounced as the dataset size exceeds 100k hours. Finally, we validate that the proposed Emilia dataset effectively supports multilingual and crosslingual speech generation, paving the way for future advancements in this field. Future work may include training effective spoof detection models to mitigate potential safety concerns arising from the use of highly spontaneous and human-like speech generation models trained with Emilia, such as the risk of synthetic spoken misinformation [44]. Additionally, further extending the Emilia dataset to the singing domain may improve singing voice generation [45].
Despite the advancements, we point out a few limitations. First, our used speaker diarization system, while SOTA, is not perfect and results in a small proportion of speech segments containing more than one speaker or overlaps, subsequently affecting speech generation performance. Integrating more advanced speaker diarization and separation techniques in the future may alleviate this issue. Second, Emilia-Pipe segments speech samples into intervals of 3 to 30 seconds. Is is observed that generating speech outside this range may sometimes lead to unexpected outcome. Adjustments in the hyper-parameters of Emilia-Pipe may be needed for specific use cases.

### v2

@@ V-C Multilingual and Crosslingual Speech Generation (EQ3)
To address EQ3, we conduct experiments to evaluate the effectiveness of Emilia for training both multilingual (where the reference and target speech are in the same language) and crosslingual (where the reference and target speech are in different languages) speech generation models.
@@
We utilize the complete Emilia-Large dataset, which encompasses six languages: English (En), Chinese (Zh), German (De), French (Fr), Japanese (Ja), and Korean (Ko), to train speech generation models. For English evaluation, we leverage Emilia-Test. The Chinese evaluation set is randomly sampled from the AISHELL-3 dataset [17]. The evaluation sets for German, French, Japanese, and Korean are sourced from Common Voice [41]. Each evaluation set contains at least 500 reference speech samples. In multilingual experiments, these reference speech samples are used to synthesize target texts in the same language. For crosslingual experiments, the reference speech samples and target texts are selected from evaluation sets of different languages.
@@
The experimental results in Table VI demonstrate the effectiveness of the Emilia-Large dataset for multilingual and crosslingual speech generation.
In multilingual generation, where the reference and target languages are the same, both AR+SoundStorm and VoiceBox achieve strong performance across all six languages. For example, AR+SoundStorm attains WERs ranging from 3.6% (Zh-Zh) to 6.3% (Ko-Ko), with S-SIM scores between 0.511 (Zh-Zh) and 0.673 (Ko-Ko), and FSD as low as 24.99 (En-En).
In crosslingual generation, where the reference and target languages differ, model performance shows moderate degradation. For instance, AR+SoundStorm’s WER increases from 6.3% (Ko-Ko) to 8.3% (En-Ko). Similarly, VoiceBox’s S-SIM decreases from 0.588 (En-En) to 0.386 (En-Zh). Nonetheless, the overall performance of both models in crosslingual generation remains competitive. These results validate the effectiveness of the Emilia-Large dataset in training strong multilingual and crosslingual speech generation models.
Furthermore, when comparing model performance to those trained exclusively on English data, the multilingual models exhibit a slight trade-off in English generation. For example, AR+SoundStorm trained on the 216k-hour multilingual Emilia-Large dataset achieves a WER of 4.9%, an S-SIM of 0.636, and an FSD of 15.31 on English evaluation samples. These metrics are marginally worse than those of its monolingual counterpart trained on a 134k-hour English-only dataset (WER=4.5%, S-SIM=0.65, FSD=14.8, as reported in Fig. 5(a)5(c)). This suggests that while multilingual data enables crosslingual capabilities, it can compromise language-specific performance. Such trade-offs are important considerations for practical applications, and future work could explore strategies to narrow these performance gaps.
@@
The Emilia-Large dataset can be effectively used to train robust multilingual and crosslingual speech generation models. However, crosslingual generation introduces moderate performance degradation, highlighting challenges in cross-language acoustic transfer. These findings underscore the value of the Emilia-Large dataset as a critical resource for advancing multilingual and crosslingual speech generation. Future work could focus on enhancing model adaptability to address crosslingual challenges.
VI Conclusion and Discussion
In conclusion, this work first introduces Emilia-Pipe, an effective and efficient preprocessing pipeline designed to transform source speech data in-the-wild into high-quality training datasets for spontaneous and human-like speech generation. Leveraging Emilia-Pipe, we construct Emilia, one of the largest open-source multilingual speech generation datasets, spanning over 101k hours across six languages, as well as its extended version, Emilia-Large, which contains 216k hours of data. Comparative analyses demonstrate that Emilia significantly outperforms traditional audio-book datasets in generating spontaneous and human-like speech. Our experiments also investigate the relationship between dataset size and speech generation performance, revealing consistent improvements with data scaling, though the trend becomes less pronounced as the dataset size exceeds 100k hours. Finally, we validate that the proposed Emilia dataset effectively supports multilingual and crosslingual speech generation, paving the way for future advancements in this field.
Future work may focus on training effective spoof detection models to address potential safety concerns associated with highly spontaneous and human-like speech generation models trained on Emilia, such as the risk of synthetic spoken misinformation [42, 43]. Additionally, expanding Emilia to include the singing/music domain could benefit singing/music voice generation [44].
Despite the advancements, we point out a few limitations. First, the speaker diarization model we use is not perfect and can result in a small proportion of speech segments containing more than one speaker or overlaps. This issue can subsequently affect speech generation performance. Integrating stronger models in the future may alleviate this issue. Second, Emilia-Pipe segments speech samples into intervals of 3 to 30 seconds. It is observed that generating speech outside this range may lead to unexpected outcomes. Adjustments in the hyper-parameters of Emilia-Pipe may be needed for specific use cases. Third, due to our resource constraints, the current Emilia-Large dataset covers only six languages. We welcome community contributions to extend the dataset to more languages with our open-source Emilia-Pipe to enhance its global applicability.